use the one mean t interval procedure to find a confidence level for the mean of the population.   x bar= 50, n=16, s=5 and the confidence level is 99%
Answer #1

x̅ = 50, s = 5, n = 16  

99% Confidence interval :  

At α = 0.01 and df = n-1 = 15, two tailed critical value, t-crit = T.INV.2T(0.01, 15) = 2.947

Lower Bound = x̅ - t-crit*s/√n = 50 - 2.947 * 5/√16 = 46.3166

Upper Bound = x̅ + t-crit*s/√n = 50 + 2.947 * 5/√16 = 53.6834

46.3166 < µ < 53.6834